OpenStudy (anonymous):

80 is 40% of what number? HELP

OpenStudy (anonymous):

(80 X 100)/40 = 200

OpenStudy (austinl):

\(\large{\dfrac{80}{x}=\dfrac{40}{100}}\) Cross multiply and divide.

OpenStudy (ivettef365):

just divide 80 / .40

OpenStudy (anonymous):

Thank you, I got it

OpenStudy (phi):

you could do this 80 is 40% of what number change the word "is" to = 80 = 40% of what number change "what number" to x 80= 40% of x change of to multiply 80 = 40% * x change 40% to 0.40 80 = 0.40 x now divide both sides by 0.4 to solve for x 80/0.4 = x